a retailer spent $48 to purchase a number of special mugs. two of them were broken in the store, but by selling each of the remaining mugs for $3 above the original cost per mug, she made a total profit of $22.
construct an equation that will allow us to solve for the # of mugs, denoted by N, that were originally purchased.(QUADRATIC EQUATIONS)
a)if the price for n mugs is $48, how can we express the cost per mug?
b)we are told that the retailer, in order to make a specific profit, sells the remaining mugs for $3 over the original cost per mug. using this imformation, and your answer to step (a), how can we express the selling price per mug?
c)since to of the mugs were broken in the store, how can we represent the # of mugs that are available for sale?
d)use your answers from step (b) and (c) to construct an expression for the amount of money made on the sale of the remaining mugs.
e)the profit on the mugs can be found by subtracting the cost of the mugs from the amount of money made on the sale of the mugs. write an equation that states this relationship between the money made, the cost, and the profit.
PLEASE HELP!!!!!!!!!!!!!!!!!!!!

Answer by Abbey(339)

Let n = the number of mugs purchased by the retailer
Let n-2 = the number of mugs available for sale (the retailer broke two)
Let (48/n) - the cost of each mug purchased by the retailer ($48 spent divided by the number of mugs)
If she sold each mug for a 3 profit
(3+(48/n)) is the sale price of each mug
(n-2)(3+(48/n)) = the total sales (the number of mugs sold*the price)
the total sales = cost + $22 profit
(n-2)(3+(48/n))=48+$22